At the end of May​, Glass ​Bottling's mixing department had​ "Total costs to account​ for" of $ 746 comma 302. Of this​ amount, $ 281 comma 772 related to direct materials​ costs, while the remainder related to conversion costs. The department had 52 comma 180 total equivalent units of direct materials and 45 comma 100 total equivalent units of conversion costs for the month. Compute the cost per equivalent unit for direct materials and the cost per equivalent unit for conversion costs.   
Answer #1

Answer

--Cost per Equivalent unit for direct material = $ 281772 / 52180 = $ 5.40

--Cost per Equivalent unit for conversion cost = $ (746302 - 281772) / 45100 units = 464530 / 45100 = $ 10.30